A liquid has a specific gravity of 0.50. What is its density in g/cm3? What is its

specific volume in cm3


/g? What is its density in lbm/ft3? What is the mass of 3.0


cm3

of this liquid? What volume is occupied by 18 g?


1
Expert's answer
2021-10-07T03:47:53-0400

density = 0.5× 1g/cm3= 0.5g/cm3

Volume = 0.5/0.5 = 1cm3